# Rate-reliability functions

### v2

@@ IV-C Extension to quantum channels
Corollary 8 and the improved bound in Eq. (21) are similarly extended to the quantum setting using the reduction from general quantum channels with an input state restriction $\mathcal{X}\subset\mathcal{S}(A)$ to cq-channels, it is clear that the rate-reliability function of a quantum channel DI $\mathcal{X}$-code is upper-bounded by Theorem 14.
IV-D Large error exponent regime
Let us start by understanding when and why is our analysis limited to the regime of small error exponents. We have to go back to the proof of Theorem 4 where, in the second inequality of Equation (12), we use the bound $-\ln F^{2}\geq 1-F$, with $F$ the fidelity between two output distributions. Notice that this bound is only a good approximation when the letterwise fidelity is close to 1. That is, when the probability distributions are similar, which can be related to a finer packing, so smaller error exponents. Conversely, large error exponents are related to wider packings, which result on fidelity values getting closer to 0, making the bound $-\ln F^{2}\geq 1-F$ grossly inaccurate. So, when applying this bound, we limit our analysis to small error exponents.
The reader could wonder now why is it at all necessary to apply such a problematic bound. Well, in order to apply our proof idea, we need a metric so that a packing can be defined, but the negative logarithm of the fidelity [second line in Equation (9)] is not a well defined distance, as it does not satisfy the triangle inequality. This is why we use the bound $-\ln F^{2}\geq 1-F$ together with the lower bound in Equation (4), conveniently obtaining an Euclidean metric. As we have already discussed, this step is a good approximation for the small exponent regime, which provides the framework for the superlinearity study and allows the recovery of the capacity results from [CDBW:DI_classical] (which were the main objectives of this work). However, it is increasingly inexact for bigger error exponents making the bound inadequate in the regime of large error exponents.
The lack of a lower bound able to well approximate the $-\ln F^{2}$ in the range of small fidelities (for which $-\ln F^{2}$ grows to infinity) and from which a well-defined distance can be extracted, suggests that our proof will probably not be trivially generalised to the large error exponent regime. Different ideas are needed. However, we can make some comments about how the rate-reliability function behaves in the large exponent regime, which are included next.
We can easily upper-bound the rate for DI over any channel $W$ more meaningfully in the regime of large error exponents starting from a transmission code for $W$. It is clear that any transmission code without randomness on the encoder and error probability $\lambda=2^{-nE_{\text{T}}}$ is also a DI code with $2^{-nE_{\text{DI}}}=\lambda_{1}=\lambda_{2}=\lambda$. Hence the rate-reliability function for transmission is a lower bound on the rate-reliability function for DI: $E_{\text{DI}}(R)\geq E_{\text{T}}(R)$. Furthermore, by using a sequential decoding of the $N=2^{nR}$ transmission messages using an $(n,N,\lambda,\lambda)$-DI code (decoding a message by trying to identify each possible output), we can bound
$2^{-nE_{\text{T}}}=\lambda\leq 2^{nR}2^{-nE_{\text{DI}}}=2^{-n(E_{\text{DI}}-R)},$
V Conclusions
We have initiated the study of the rate-reliability tradeoff between the message length and the exponents of the two error probabilities occurring in DI. The main finding is that, unlike in the communication problem, imposing exponentially (in the block length $n$) small errors of first and second kind makes the slightly superlinear scaling of the message length disappear, and replaces it with a “regular” linear-scale rate. We do not have exact formulas for the rate-reliability function, whose derivation remains as the main open question for further research. But our results are strong enough to show that for sufficiently small error exponents a universal behavior manifests itself: the rate is essentially proportional to the Minkowski dimension of the set $\sqrt{\!\widetilde{\mathcal{X}}}$ (up to a factor between $\frac{1}{4}$ and $\frac{1}{2}$) multiplied by the logarithm of the inverse exponent.
Technically, we derive upper and lower bounds on the DI rate for fixed reliability exponents and at finite block length. While the proofs are mainly inspired by those in [CDBW:DI_classical], the analysis in the non-asymptotic regime allows us to separate the information theoretic from the geometric aspects of the problem, resulting in the bounding of the DI rate in terms of error-exponent-dependent packings (for the lower bound, Theorem 4) and coverings (for the upper bound, Theorem 7). The geometric inspection of the bounds in the regime of small errors yields the appearance of the lower and upper Minkowski dimensions featuring in the main results of [CDBW:DI_classical]. There, these two aspects appeared intertwined, as the covering and packing radii were made to directly depend on the block length (which is treated asymptotically), perhaps obscuring the main ideas behind the proof. We also show how to approach the special case of channels with a vanishing Minkowski dimension of the output probability set through a couple of examples, and extend the reliability results to classical-quantum channels and quantum channels under some input restriction. We approach the case of big error exponents for which our lower bound becomes trivial, and comment on some characteristics that the reliability curve must have in that regime, leaving a more thorough study for future work. These sections complete our first reliability study of deterministic identification making it a parallel to the capacity study [CDBW:DI_classical], in the sense that every capacity result there can now be mapped to a reliability result here, in this paper.
The present results explain somewhat the origin of the surprising superlinear performance characteristic of deterministic identification: here we see that it has its roots in a sufficiently slow convergence of the errors to zero. By decoupling the error exponents from the block length, we discover that they determine the geometric scale at which the packing and covering properties of $\sqrt{\!\widetilde{\mathcal{X}}}$ determine the (linear) DI rate. Thus, the Minkowski dimension emerges in the regime of small error exponents. It remains as an open question to determine the DI code scaling when merely one of the two errors is exponentially small, i.e. $E_{1}>0,\,E_{2}=0$ or $E_{1}=0,\,E_{2}>0$. This is an interesting case as it corresponds to the Stein’s setting of asymmetric hypothesis testing: the alternative hypothesis, if correct, should be detected with decent probability; but the null hypothesis should be falsely rejected only with exponentially small probability (or the other way round). In this scenario, our converse gives the previous upper bound of $\frac{1}{2}d_{M}\,n\log n$ on the message length, but our current code constructions only seem to attain linear scaling.

### v4

@@ VI Extension to quantum channels
Corollary 8 and the improved bound in Equation (41) are similarly extended to the quantum setting using again the reduction from general quantum channels with an input state restriction $\mathcal{X}\subset\mathcal{S}(A)$ to cq-channels.
@@
$nE_{2,1}$ $\leq D_{h}^{\lambda_{1,2}}(W_{u_{j}}\|W_{u_{k}})$
$\leq\widetilde{D}_{\alpha}(W_{u_{j}}\|W_{u_{k}})-\frac{\alpha}{\alpha-1}\log(1-\lambda_{1,2})$
$\leq\sum_{i=1}^{n}\widetilde{D}_{\alpha}(W_{x_{i}}\|W_{x_{i}^{\prime}})-\frac{\alpha}{\alpha-1}\log(1-\lambda_{1,2}),$ (94)
where
$\widetilde{D}_{\alpha}(\rho\|\sigma)=\frac{1}{\alpha-1}\log\operatorname{Tr}\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}$ (95)
is the sandwiched Rényi relative entropy. Next, we use the following continuity bound for the latter, leveraging the eigenvalue lower bound:
Lemma 19 (Cf. Bluhm et al. [40, Cor. 5.7]).
[…]
$\widetilde{D}_{\alpha}(\rho\|\sigma)$ $\leq\log(1\!+\!t)+\frac{1}{\alpha-1}\log\!\left(\!1\!+\!t\omega^{1-\alpha}\!+\!\frac{t^{\alpha}}{(1+t)^{\alpha-1}}\!\right)$
$\leq\frac{t}{\ln 2}\left[1+\frac{1}{\alpha-1}\left(\frac{1}{\omega^{\alpha-1}}-\left(\frac{t}{1+t}\right)^{\alpha-1}\right)\right]$
$\leq t\,K(\omega,\alpha).$ (96)
Using this, we can convert the above Equation (VI) into bounds on sums of trace distances:
$\sum_{i=1}^{n}\frac{1}{2}\|W_{x_{i}}-W_{x_{i}^{\prime}}\|_{1}\geq\begin{cases}\frac{E_{2}}{K(\omega,\alpha)}n+\frac{\alpha\log(1-\lambda_{1})}{(\alpha-1)K(\omega,\alpha)},\\[8.61108pt] \frac{E_{1}}{K(\omega,\alpha)}n+\frac{\alpha\log(1-\lambda_{2})}{(\alpha-1)K(\omega,\alpha)}.\end{cases}$ (97)
Either right hand side has the form $\gamma n-\Delta$. Now, partition $\widetilde{\mathcal{X}}=\bigcup_{\ell=1}^{L}\widetilde{\mathcal{X}}_{\ell}$ such that the diameter of each $\widetilde{\mathcal{X}}_{\ell}$ (with respect to the trace distance) is $\leq\gamma/2$ – this could for example come from a $\gamma/4$-covering, which has cardinality $L=\Gamma_{\gamma/4}(\widetilde{\mathcal{X}})$. This means that for any $x^{n},\,{x^{\prime}}^{n}$ such that $W_{x^{n}},\,W_{{x^{\prime}}^{n}}\in\widetilde{\mathcal{X}}_{\ell^{n}}=\widetilde{\mathcal{X}}_{\ell_{1}}\times\cdots\times\widetilde{\mathcal{X}}_{\ell_{n}}$ (that is, from the same region of the $n$-letter partition), $\sum_{i=1}^{n}\frac{1}{2}\|W_{x_{i}}-W_{x_{i}^{\prime}}\|_{1}\leq\frac{\gamma}{2}n,$ which is $<\gamma n-\Delta$ for all sufficiently large $n$. In other words, each $\widetilde{\mathcal{X}}_{\ell^{n}}$ can contain at most one $W_{u_{j}}$, thus $N\leq L^{n}$, or $R(n)\leq\frac{1}{n}\log N\leq\log L$. Note that $L$ is a function of $\gamma$, which in turn is determined by $E_{1,2}$, $\alpha>1$ and $\omega>0$, but in any case a constant with respect to $n$. Thus we proved the following quantum generalisation of Theorem 14:
Theorem 20.
[…]
On the other hand, without this eigenvalue lower bound there are examples of cq-channels attaining linearithmic DI rates in the Sanov setting with $\lambda_{1}=0$ (i.e. $E_{1}=+\infty$) and $\lambda_{2}\rightarrow 0$ [38, Sect. V]. Crucially, the codes constructed there are non-simultaneous. We do not know if linearithmic rates are possible for cq-channels in the Stein setting.
VII Conclusions
We have initiated the study of the rate-reliability tradeoff between the message length and the exponents of the two error probabilities occurring in DI. The main finding is that, unlike in the communication problem, imposing exponentially (in the block length $n$) small errors of first and second kind makes the slightly superlinear scaling of the message length disappear, and replaces it with a “regular” linear-scale rate. Even in the case where only one of the two errors vanishes exponentially, superlinearity is lost. We do not have exact formulas for the rate-reliability function, whose derivation remains as the main open question for further research. But our results are strong enough to show that for sufficiently small error exponents a universal behaviour manifests itself: the rate is essentially proportional to the Minkowski dimension of the set $\sqrt{\!\widetilde{\mathcal{X}}}$ (up to a factor between $\frac{1}{4}$ and $\frac{1}{2}$) multiplied by the logarithm of the inverse exponent.
Technically, we derive upper and lower bounds on the DI rate for fixed reliability exponents and at finite block length. While the proofs are mainly inspired by those in [17], the analysis in the non-asymptotic regime allows us to separate the information theoretic from the geometric aspects of the problem, resulting in the bounding of the DI rate in terms of error-exponent-dependent packings (for the lower bound, Theorem 4) and coverings (for the upper bound, Theorem 7). The geometric inspection of the bounds in the regime of small errors yields the appearance of the lower and upper Minkowski dimensions featuring in the main results of [17]. There, these two aspects appeared intertwined, as the covering and packing radii were made to directly depend on the block length (which is treated asymptotically), perhaps obscuring the main ideas behind the proof. The finite block length approach started in this paper brings this still very abstract but exciting theory closer to potential implementation.
The present results go some way towards explaining the origin of the surprising superlinear performance characteristic of deterministic identification: here we see that it has its roots in a sufficiently slow convergence of both errors to zero. By decoupling the error exponents from the block length, we discover that they determine the geometric scale at which the packing and covering properties of $\sqrt{\!\widetilde{\mathcal{X}}}$ determine the (linear) DI rate. Thus, the Minkowski dimension emerges in the regime of small error exponents.
We furthermore show how to approach the special case of channels with a vanishing Minkowski dimension of the output probability set through a couple of examples, and extend the reliability results to classical-quantum channels and quantum channels under some input restriction. Thirdly, we approach the case of large error exponents for which our lower bound becomes trivial, and comment on some characteristics that the reliability curve must have in that regime, leaving a more thorough study for future work.
While the results up to that point make sense mainly in the regime of positive and comparable error exponents, we go beyond this setting in Section V, where the regime of extremely differently behaving errors is investigated: one error vanishes exponentially fast while the other is just bounded by a constant. Also in this case (which is closely related to the settings of Stein’s Lemma and Sanov’s Theorem in Hypothesis testing) we find that the superlinear scaling of DI codes is lost. These sections complete our first reliability study of deterministic identification making it a parallel to the capacity study [17], in the sense that every capacity result there can now be obtained through the reliability results in this paper, in particular error regimes.
While our analysis is based on entropy typicality arguments, which are more naturally extendable to continuous and quantum settings, we also demonstrate that it recovers known results previously derived using the method of types and strong typicality [9] (see Example 12). It is well-known that in classical channel coding over discrete memoryless channels the method of types can be used to derive optimal error exponents. But such techniques are not applicable in our more general setting, which includes continuous and quantum channels where empirical distributions (types) are not well defined. We think that the non-matching upper and lower bounds on the DI capacity cannot simply be attributed to the use of the “wrong” typicality concept.
As already mentioned, the biggest open problem is closing the gap between our upper and lower bounds on the rate-reliability functions and capacity results. Indeed, our bounds could be taken to suggest that the linearithmic capacity of DI over general channel is $\dot{C}_{\text{DI}}(W)=\alpha\underline{d}_{M}\!\left(\!\!\sqrt{\!\widetilde{\mathcal{X}}}\right)$, with $\alpha\in[\frac{1}{4},\frac{1}{2}]$ a constant which we still need to fix. The recent result that for Gaussian channels $\alpha\geq\frac{3}{8}$ [41], together with the difficulty of lowering the upper bound, might be taken to indicate that our code construction (packing with hypothesis testing lemma), and hence the achievability bound, can be improved. Furthermore, we have examples of particular channels (both classical [17, Sect. V-F] and quantum [38]) for which suitable decoding sets beyond the entropy typical set can be devised which attain the capacity upper bound.